OpenStudy (anonymous):

Which of the following is a polynomial with roots 2, -2i, and 2i ? Choose one answer. A. x3 - 4x2 + 4x - 16 B. x3 + 4x2 + 4x + 16 C. x3 + 2x2 + 4x + 8 D. x3 - 2x2 + 4x - 8

OpenStudy (anonymous):

the one that looks like \[(x-2)(x^2+4)\] after you multiply out

Nnesha (nnesha):

do you have calculator ???

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

you don't need one, just multiply btw of course you have a calculator! you are on a computer

Nnesha (nnesha):

no if you have calculator you can get half help from it and yeah good point about computer cal. but i asking ti84 one real in hands

OpenStudy (anonymous):

i picked c, is that correct?

OpenStudy (anonymous):

c is usually correct is it \[x^3-2 x^2+4 x-8\]?

Nnesha (nnesha):

yes that is correct

OpenStudy (anonymous):

thank you! could you guys help me with this problem? Which of the following is a polynomial with roots - square root of 5, square root of 5, and -3 ? Choose one answer. A. x3 - 2x2 - 3x + 6 B. x3 + 2x2 - 3x - 6 C. x3 - 3x2 - 5x + 15 D. x3 + 3x2 - 5x - 15

OpenStudy (anonymous):

\[(x+3)(x-\sqrt5)(x+\sqrt5)=(x+3)(x^2-5)\]

OpenStudy (anonymous):

thank you :) this is the last question! What are the possible number of positive, negative, and complex zeros of f(x) = -x6 + x5- x4 + 4x3 - 12x2 + 12 ? Choose one answer. A. Positive: 4, 2, or 0; Negative: 2 or 0; Complex: 6, 4, 2, or 0 B. Positive: 3 or 1; Negative: 3 or 1; Complex: 4, 2, or 0 C. Positive: 5, 3, or 1; Negative: 1; Complex: 4, 2, or 0 D. Positive: 2 or 0; Negative: 2 or 0; Complex: 6, 4, or 2

Nnesha (nnesha):

do you know how to factor polynomial ???

OpenStudy (anonymous):

no i dont. im really bad at math and i take online classes so we just get problems with no examples on how to do things :/

Nnesha (nnesha):

so you should learn this i don't know if you got above two for last one what is your highest degreee

Nnesha (nnesha):

do you know about degree ??? :)

Nnesha (nnesha):

 exponent known as degree so what is highest degree